# Analyzing Cognitive Differences Among Large Language Models through the Lens of Social Worldview

### v1

@@ Appendix C Automated Multi-Agent Prompting Framework
Taxonomy Alignment Agent (Prompt J)
@@
Questions scoring below the threshold (score $<3$) are explicitly flagged for refinement, with brief justifications documenting alignment assessments.
Semantic Validation Agent (Prompt J)
@@
Refinement Agent (Prompt J)
@@
Appendix D Questionnaire Description and Evaluation
D.1 Description
@@ D.2 Evaluation Methodology
Confirmatory Factor Analysis (CFA)
Confirmatory Factor Analysis tests how effectively our questionnaire items align with the proposed theoretical structure of the SWT. CFA assumes a factor model of the form:
$\mathbf{x}=\mathbf{\Lambda}\mathbf{\xi}+\mathbf{\delta}$
where $\mathbf{x}$ is a vector of observed questionnaire item responses, $\mathbf{\Lambda}$ is the factor loading matrix representing the relationships between observed items and latent SWT dimensions, $\mathbf{\xi}$ denotes latent factors (dimensions), and $\mathbf{\delta}$ is the vector of measurement errors. We evaluate the CFA model fit using several standard indices:
Comparative Fit Index (CFI)
CFI measures the improvement in fit of the specified CFA model compared to a baseline independence model:
$\text{CFI}=1-\frac{\chi^{2}_{\text{model}}-df_{\text{model}}}{\chi^{2}_{\text{baseline}}-df_{\text{baseline}}}$
CFI values closer to $1.0$ (generally $\geq 0.90$) indicate excellent model fit.
Tucker–Lewis Index (TLI)
TLI is a variant of CFI adjusted for model complexity:
$\text{TLI}=\frac{\chi^{2}_{\text{baseline}}/df_{\text{baseline}}-\chi^{2}_{\text{model}}/df_{\text{model}}}{\chi^{2}_{\text{baseline}}/df_{\text{baseline}}-1}$
TLI values equal to or greater than $0.90$ indicate an acceptable fit to the data.
Root Mean Square Error of Approximation (RMSEA)
RMSEA evaluates how well the specified model approximates reality, penalizing for complexity:
$\text{RMSEA}=\sqrt{\frac{\chi^{2}_{\text{model}}-df_{\text{model}}}{df_{\text{model}}(N-1)}}$
where $N$ represents sample size. RMSEA values less than or equal to $0.08$ indicate an acceptable approximation.
Standardized Root Mean Square Residual (SRMR)
SRMR quantifies the average discrepancy between observed and model-predicted correlations:
$\text{SRMR}=\sqrt{\frac{2\sum_{i=1}^{p}\sum_{j<i}^{p}(r_{ij}-\hat{r}_{ij})^{2}}{p(p-1)}}$
where $r_{ij}$ and $\hat{r}_{ij}$ are observed and predicted correlations between items $i$ and $j$, respectively, and $p$ is the number of items. SRMR values less than or equal to $0.08$ reflect a good model fit.
D.3 Manual Evaluation
To validate the quality of our automatically generated questionnaire items, we conducted a manual evaluation. Expert annotators independently assessed all generated questions across each SWT dimension. Each question was evaluated using binary scoring (1 = acceptable, 0 = unacceptable) along two criteria: (1) Dimension Alignment, defined as whether the question correctly aligns with the intended SWT dimension; and (2) Clarity, indicating whether the question is clear, unambiguous, and effectively measurable via the intended Likert scale. Accuracy scores (%) were then computed for each sub-dimension based on these binary assessments, as summarized in Table 4.
@@ D.3 Manual Evaluation
Appendix E Additional Descriptive Results
Figure 4: Response mean value for main models under each condition
Figure 5: Response distribution for main models under each condition
Dimension llama3.3-70b qwen2.5-72b deepseek-v3 gemini2.0-flash gemma3-27b glm4-9b gpt-4o internlm2.5-20b phi4
Egalitarianism 4.51 [4.38, 4.64] 4.23 [4.12, 4.34] 4.56 [4.44, 4.69] 4.41 [4.26, 4.55] 4.42 [4.29, 4.56] 4.34 [4.23, 4.45] 4.53 [4.40, 4.65] 4.22 [4.11, 4.33] 4.47 [4.35, 4.59]
Fatalism 2.92 [2.76, 3.09] 2.88 [2.78, 2.97] 2.83 [2.69, 2.98] 2.57 [2.41, 2.72] 2.81 [2.66, 2.96] 2.91 [2.81, 3.00] 2.94 [2.79, 3.08] 3.06 [3.01, 3.12] 3.09 [2.99, 3.19]
Hierarchy 3.64 [3.46, 3.83] 3.46 [3.32, 3.59] 3.62 [3.45, 3.80] 3.44 [3.25, 3.64] 3.44 [3.25, 3.63] 3.66 [3.53, 3.79] 3.69 [3.52, 3.87] 3.53 [3.43, 3.64] 3.59 [3.42, 3.76]
Individualism 4.21 [4.07, 4.35] 3.96 [3.86, 4.07] 4.19 [4.08, 4.30] 4.21 [4.10, 4.33] 4.04 [3.90, 4.17] 3.91 [3.80, 4.02] 4.13 [4.01, 4.25] 3.90 [3.80, 4.00] 4.04 [3.93, 4.15]
@@ Appendix E Additional Descriptive Results
Dimension llama3.3-70b qwen2.5-72b deepseek-v3 gemini2.0-flash gemma3-27b glm4-9b gpt-4o internlm2.5-20b phi4
Egalitarianism 4.41 [4.27, 4.55$\downarrow$ 4.28 [4.17, 4.38$\uparrow$ 4.54 [4.41, 4.67] 4.44 [4.29, 4.58] 4.38 [4.24, 4.52$\downarrow$ 4.30 [4.19, 4.41] 4.52 [4.40, 4.64] 4.27 [4.16, 4.38$\uparrow$ 4.53 [4.41, 4.65$\uparrow$
Fatalism 2.76 [2.58, 2.94$\downarrow$ 2.81 [2.71, 2.91$\downarrow$ 2.74 [2.60, 2.89$\downarrow$ 2.56 [2.39, 2.72] 2.77 [2.64, 2.91] 3.00 [2.92, 3.08$\uparrow$ 2.87 [2.72, 3.02] 3.08 [3.03, 3.14] 3.07 [2.96, 3.18$\uparrow$
Hierarchy 3.56 [3.38, 3.75$\downarrow$ 3.44 [3.31, 3.58] 3.58 [3.40, 3.76] 3.46 [3.26, 3.67] 3.46 [3.28, 3.63] 3.66 [3.54, 3.78] 3.64 [3.47, 3.82] 3.58 [3.47, 3.68$\uparrow$ 3.66 [3.48, 3.83$\uparrow$
Individualism 4.03 [3.87, 4.18$\downarrow$ 3.99 [3.89, 4.10] 4.16 [4.05, 4.28] 4.24 [4.12, 4.36] 4.06 [3.94, 4.19] 3.88 [3.78, 3.98] 4.16 [4.05, 4.27] 3.94 [3.84, 4.04] 4.10 [3.99, 4.21$\uparrow$
@@
Feedback Dimension llama3.3-70b qwen2.5-72b deepseek-v3 gemini2.0-flash gemma3-27b glm4-9b gpt-4o internlm2.5-20b phi4
None Egalitarianism 4.39 [4.23, 4.55] 4.26 [4.14, 4.37] 4.50 [4.37, 4.63] 3.99 [3.86, 4.13$\downarrow$ 3.62 [3.48, 3.76$\downarrow$ 4.71 [4.62, 4.79$\uparrow$ 4.38 [4.26, 4.51$\downarrow$ 4.04 [3.91, 4.18$\downarrow$ 4.59 [4.49, 4.70$\uparrow$
Fatalism 2.66 [2.46, 2.86$\downarrow$ 2.92 [2.83, 3.01$\uparrow$ 2.83 [2.69, 2.97$\uparrow$ 2.86 [2.73, 2.99$\uparrow$ 2.66 [2.51, 2.80] 3.71 [3.60, 3.83$\uparrow$ 2.85 [2.71, 2.99] 3.34 [3.23, 3.44$\uparrow$ 3.33 [3.19, 3.46$\uparrow$
Hierarchy 3.41 [3.20, 3.63$\downarrow$ 3.42 [3.29, 3.56] 3.55 [3.37, 3.73] 3.03 [2.89, 3.18$\downarrow$ 2.83 [2.69, 2.97$\downarrow$ 3.98 [3.85, 4.10$\uparrow$ 3.38 [3.21, 3.54$\downarrow$ 3.24 [3.14, 3.34$\downarrow$ 3.72 [3.54, 3.89$\uparrow$
Individualism 4.01 [3.84, 4.17] 3.98 [3.87, 4.09] 4.03 [3.90, 4.16$\downarrow$ 3.64 [3.52, 3.76$\downarrow$ 2.98 [2.83, 3.12$\downarrow$ 4.17 [4.08, 4.26$\uparrow$ 3.83 [3.71, 3.96$\downarrow$ 3.58 [3.46, 3.69$\downarrow$ 4.22 [4.12, 4.32$\uparrow$
Little Egalitarianism 4.43 [4.28, 4.58] 4.28 [4.17, 4.38] 4.53 [4.40, 4.65] 4.19 [4.06, 4.32$\downarrow$ 3.91 [3.76, 4.05$\downarrow$ 4.73 [4.65, 4.81$\uparrow$ 4.47 [4.35, 4.59$\downarrow$ 4.16 [4.03, 4.28$\downarrow$ 4.61 [4.50, 4.71$\uparrow$
Fatalism 2.71 [2.52, 2.90$\downarrow$ 2.98 [2.90, 3.07$\uparrow$ 2.88 [2.74, 3.02$\uparrow$ 2.90 [2.75, 3.05$\uparrow$ 2.90 [2.75, 3.05$\uparrow$ 3.77 [3.67, 3.88$\uparrow$ 2.86 [2.71, 3.00] 3.54 [3.45, 3.63$\uparrow$ 3.31 [3.18, 3.45$\uparrow$
Hierarchy 3.51 [3.31, 3.71$\downarrow$ 3.46 [3.32, 3.59] 3.57 [3.39, 3.75] 3.20 [3.05, 3.35$\downarrow$ 3.03 [2.88, 3.18$\downarrow$ 4.03 [3.91, 4.15$\uparrow$ 3.54 [3.37, 3.71$\downarrow$ 3.46 [3.35, 3.56$\downarrow$ 3.70 [3.53, 3.87$\uparrow$
Individualism 4.02 [3.86, 4.18] 3.99 [3.89, 4.10] 4.13 [4.01, 4.25] 3.83 [3.71, 3.95$\downarrow$ 3.31 [3.17, 3.45$\downarrow$ 4.28 [4.18, 4.37$\uparrow$ 4.06 [3.95, 4.17$\downarrow$ 3.68 [3.57, 3.80$\downarrow$ 4.22 [4.12, 4.32$\uparrow$
Most Egalitarianism 4.58 [4.43, 4.73$\uparrow$ 4.28 [4.17, 4.38] 4.55 [4.42, 4.68] 4.46 [4.32, 4.60] 4.51 [4.37, 4.64$\uparrow$ 4.76 [4.68, 4.83$\uparrow$ 4.52 [4.40, 4.64] 4.41 [4.33, 4.48$\uparrow$ 4.67 [4.56, 4.77$\uparrow$
Fatalism 2.78 [2.57, 2.99] 2.89 [2.80, 2.98$\uparrow$ 2.94 [2.78, 3.11$\uparrow$ 2.66 [2.49, 2.84$\uparrow$ 2.91 [2.75, 3.06$\uparrow$ 3.84 [3.74, 3.93$\uparrow$ 2.89 [2.74, 3.05$\uparrow$ 3.91 [3.87, 3.96$\uparrow$ 3.20 [3.07, 3.33$\uparrow$
Hierarchy 3.59 [3.38, 3.81] 3.44 [3.31, 3.58] 3.64 [3.47, 3.82$\uparrow$ 3.58 [3.38, 3.77$\uparrow$ 3.51 [3.33, 3.70$\uparrow$ 4.19 [4.09, 4.29$\uparrow$ 3.64 [3.47, 3.82] 3.98 [3.91, 4.05$\uparrow$ 3.73 [3.55, 3.90$\uparrow$
Individualism 4.11 [3.94, 4.27$\uparrow$ 3.99 [3.89, 4.10] 4.24 [4.14, 4.35$\uparrow$ 4.26 [4.15, 4.38] 4.20 [4.06, 4.34$\uparrow$ 4.45 [4.37, 4.53$\uparrow$ 4.16 [4.05, 4.27] 4.15 [4.08, 4.22$\uparrow$ 4.23 [4.13, 4.33$\uparrow$
Table 7: Responses of Mean Value and 95% Confidence Interval across Feedback Conditions. $\uparrow$: significant higher than self-awareness; $\downarrow$: significant lower than self-awareness, by paired t-test
@@
Notably, Table 7 summarizes the mean Likert-scale responses and 95% confidence intervals across three feedback conditions—None, Little, and Most—indicating how many of five peers agreed with the model’s prior answer. This manipulation probes the responsiveness of large language models (LLMs) to external social feedback, revealing how consensus cues shape value expression.
Table 7 presents the effects of feedback-based prompting—ranging from full disagreement (None), partial agreement (Little), to strong consensus (Most)—on the expression of sociopolitical values across nine LLMs. Each condition is compared against the self-awareness responses via paired t-tests. While self-awareness introduces a social-referencing framing, feedback conditions simulate varying levels of external agreement, which may elicit either value reinforcement or moderation.
@@
Figure 6: Base: Responses Correlation across Different Models
Figure 7: Self-Awareness: Responses Correlation across Different Models
Figure 8: Self-Awareness: Responses Correlation across Different Models under Different Feedback Settings
Appendix F Details for Baseline Worldview Profiles Analysis (RQ1)
F.1 Deriving LLM Personas
@@
F.1.1 Dimensionality Reduction via Structural Equation Modeling
@@
F.1.2 Persona Identification via Latent Profile Analysis
@@
F.1.3 Persona Interpretation and Visualization
@@
Dimension $\text{Dimension salience}(k,d)=\begin{cases}\text{High (\textuparrow)},&\mu_{kd}\geq\tau\\[4.0pt] \text{Low (\textdownarrow)},&\mu_{kd}\leq-\tau\\[4.0pt] \text{Neutral},&|\mu_{kd}|<\tau\end{cases}$
@@ F.1.4 Results
Table 8: Persona clusters, narrative labels, concise “vibes”, and associated LLMs (basic condition).
@@
Qwen2.5-72B-Instruct $-0.03$ $0.24$ $-0.08$ $0.10$ 2
deepseek-chat $0.07$ $-0.16$ $0.06$ $-0.08$ 3
gemini-2.0-flash $0.16$ $-0.09$ $0.10$ $-0.15$ 5
gemma-3-27B-it $-0.09$ $-0.14$ $0.02$ $-0.14$ 0
GLM-4-9B-chat $-0.09$ $0.16$ $-0.10$ $0.13$ 2
GPT-4o $0.14$ $-0.11$ $0.07$ $-0.13$ 5
InternLM-2.5-20B-chat $-0.28$ $0.25$ $-0.14$ $0.21$ 1
Llama-3.3-70B-instruct $0.13$ $-0.12$ $0.08$ $-0.05$ 3
Phi-4 $-0.01$ $-0.01$ $-0.01$ $0.10$ 4
Table 9: Latent worldview scores (z-standardised) and assigned persona cluster for each LLM under the baseline condition.
@@
0 $-0.090$ $-0.139$ $0.023$ $-0.138$
1 $-0.282$ $0.247$ $-0.141$ $0.207$
2 $-0.060$ $0.197$ $-0.092$ $0.115$
3 $0.099$ $-0.143$ $0.072$ $-0.064$
4 $-0.010$ $-0.013$ $-0.007$ $0.102$
5 $0.152$ $-0.102$ $0.082$ $-0.136$
Table 10: Cluster centroids (z-standardised) for each persona across the four latent worldview dimensions. Positive values indicate above-average alignment with that dimension, negative values indicate below-average alignment.
Our latent persona analysis revealed six distinct cognitive personas across the nine LLMs examined. Each persona represents a coherent worldview characterized by systematic patterns of alignment along four dimensions: Hierarchy, Egalitarianism, Individualism, and Fatalism. The detailed results, including model-level scores, cluster centroids, and concise narrative descriptions, are summarized in Tables 89, and 10. We narrate these personas below to illustrate distinct patterns in LLM cognitive profiles.

### v2

@@ Appendix A Automated Multi-Agent Prompting Framework
Taxonomy Alignment Agent (Prompt H)
@@
Questions scoring below the threshold (score <3<3) are explicitly flagged for refinement, with brief justifications documenting alignment assessments.
Semantic Validation Agent (Prompt H)
@@
Refinement Agent (Prompt H)
@@
Appendix B Questionnaire Description and Evaluation
B.1 Description
@@ B.2 Evaluation Methodology
B.3 Manual Evaluation
To ensure the quality and conceptual validity of the SWQ dataset, we conducted a comprehensive manual evaluation. Two domain experts, both researchers with peer-reviewed publications in computational social science and specific expertise in Cultural Theory, independently assessed all generated items following a strict annotation protocol.
Each item was evaluated using binary scoring (1 = acceptable, 0 = unacceptable) along two critical dimensions: (1) Dimension Alignment, verifying that the item accurately operationalizes the intended SWT construct rather than mere surface-level phrasing; and (2) Clarity, ensuring the item is unambiguous and effectively measurable via the intended Likert scale. To guarantee the reliability of the human annotation, we computed the inter-annotator agreement, yielding a strong Cohen’s Kappa ($\kappa=0.86$, $p<0.01$). This high level of agreement confirms robust consistency between experts. Aggregated accuracy scores based on these expert assessments are summarized in Table 3.
@@ B.3 Manual Evaluation
Appendix C Additional Descriptive Results
@@
Model Egalitarianism Fatalism Hierarchy Individualism
llama3.3-70b 4.51 [4.38, 4.64] 2.92 [2.76, 3.09] 3.64 [3.46, 3.83] 4.21 [4.07, 4.35]
qwen2.5-72b 4.23 [4.12, 4.34] 2.88 [2.78, 2.97] 3.46 [3.32, 3.59] 3.96 [3.86, 4.07]
deepseek-v3 4.56 [4.44, 4.69] 2.83 [2.69, 2.98] 3.62 [3.45, 3.80] 4.19 [4.08, 4.30]
gemini2.0-flash 4.41 [4.26, 4.55] 2.57 [2.41, 2.72] 3.44 [3.25, 3.64] 4.21 [4.10, 4.33]
gemma3-27b 4.42 [4.29, 4.56] 2.81 [2.66, 2.96] 3.44 [3.25, 3.63] 4.04 [3.90, 4.17]
glm4-9b 4.34 [4.23, 4.45] 2.91 [2.81, 3.00] 3.66 [3.53, 3.79] 3.91 [3.80, 4.02]
gpt-4o 4.53 [4.40, 4.65] 2.94 [2.79, 3.08] 3.69 [3.52, 3.87] 4.13 [4.01, 4.25]
internlm2.5-20b 4.22 [4.11, 4.33] 3.06 [3.01, 3.12] 3.53 [3.43, 3.64] 3.90 [3.80, 4.00]
phi4 4.47 [4.35, 4.59] 3.09 [2.99, 3.19] 3.59 [3.42, 3.76] 4.04 [3.93, 4.15]
@@
Model Egalitarianism Fatalism Hierarchy Individualism
[…]
gemini2.0-flash 4.44 [4.29, 4.58] 2.56 [2.39, 2.72] 3.46 [3.26, 3.67] 4.24 [4.12, 4.36]
[…]
gpt-4o 4.52 [4.40, 4.64] 2.87 [2.72, 3.02] 3.64 [3.47, 3.82] 4.16 [4.05, 4.27]
[…]
Table 6: Responses of Mean Value and 95% Confidence Interval under Feedback = None. $\uparrow$: significant higher than self-awareness; $\downarrow$: significant lower than self-awareness, by paired t-test
Model Egalitarianism Fatalism Hierarchy Individualism
[…]
Table 7: Responses of Mean Value and 95% Confidence Interval under Feedback = Little. $\uparrow$: significant higher than self-awareness; $\downarrow$: significant lower than self-awareness, by paired t-test
Model Egalitarianism Fatalism Hierarchy Individualism
[…]
Table 8: Responses of Mean Value and 95% Confidence Interval under Feedback = Most. $\uparrow$: significant higher than self-awareness; $\downarrow$: significant lower than self-awareness, by paired t-test
Model Egalitarianism Fatalism Hierarchy Individualism
[…]
@@
Notably, Tables 6, 7, and 8 summarize the mean Likert-scale responses and 95% confidence intervals across three feedback conditions—None, Little, and Most—indicating how many of five peers agreed with the model’s prior answer. This manipulation probes the responsiveness of large language models (LLMs) to external social feedback, revealing how consensus cues shape value expression.
Tables 6, 7, and 8 present the effects of feedback-based prompting—ranging from full disagreement (None), partial agreement (Little), to strong consensus (Most)—on the expression of sociopolitical values across nine LLMs. Each condition is compared against the self-awareness responses via paired t-tests. While self-awareness introduces a social-referencing framing, feedback conditions simulate varying levels of external agreement, which may elicit either value reinforcement or moderation.
@@
Appendix D Details for Baseline Worldview Profiles Analysis (RQ1)
D.1 Deriving LLM Personas
@@
D.1.1 Dimensionality Reduction via Structural Equation Modeling
@@
D.1.2 Persona Identification via Latent Profile Analysis
@@
D.1.3 Persona Interpretation and Visualization
@@
$\text{Dimension salience}(k,d)=\begin{cases}\text{High (\textuparrow)},&\mu_{kd}\geq\tau\\[4.0pt] \text{Low (\textdownarrow)},&\mu_{kd}\leq-\tau\\[4.0pt] \text{Neutral},&|\mu_{kd}|<\tau\end{cases}$
@@ D.1.4 Results
Table 10: Latent worldview scores (z‑standardised) and assigned persona cluster for each LLM under the baseline condition.
@@
Qwen2.5‑72B‑Instruct 0.03-0.03 $0.24$ 0.08-0.08 $0.10$ 2
deepseek‑chat $0.07$ 0.16-0.16 $0.06$ 0.08-0.08 3
gemini‑2.0‑flash $0.16$ 0.09-0.09 $0.10$ 0.15-0.15 5
gemma‑3‑27B‑it 0.09-0.09 0.14-0.14 $0.02$ 0.14-0.14 0
GLM‑4‑9B‑chat 0.09-0.09 $0.16$ 0.10-0.10 $0.13$ 2
GPT‑4o $0.14$ 0.11-0.11 $0.07$ 0.13-0.13 5
InternLM‑2.5‑20B‑chat 0.28-0.28 $0.25$ 0.14-0.14 $0.21$ 1
Llama‑3.3‑70B‑instruct $0.13$ 0.12-0.12 $0.08$ 0.05-0.05 3
Phi‑4 0.01-0.01 0.01-0.01 0.01-0.01 $0.10$ 4
Table 11: Cluster centroids (z‑standardised) for each persona across the four latent worldview dimensions. Positive values indicate above‑average alignment with that dimension, negative values indicate below‑average alignment.
@@
0 0.090-0.090 0.139-0.139 $0.023$ 0.138-0.138
1 0.282-0.282 $0.247$ 0.141-0.141 $0.207$
2 0.060-0.060 $0.197$ 0.092-0.092 $0.115$
3 $0.099$ 0.143-0.143 $0.072$ 0.064-0.064
4 0.010-0.010 0.013-0.013 0.007-0.007 $0.102$
5 $0.152$ 0.102-0.102 $0.082$ 0.136-0.136
Our latent persona analysis revealed six distinct cognitive personas across the nine LLMs examined. Each persona represents a coherent worldview characterized by systematic patterns of alignment along four dimensions: Hierarchy, Egalitarianism, Individualism, and Fatalism. The detailed results, including model-level scores, cluster centroids, and concise narrative descriptions, are summarized in Tables 910, and 11. We narrate these personas below to illustrate distinct patterns in LLM cognitive profiles.